PROBLEM SOLVING STRATEGY 221 Electric forces and Coulombs law MODEL: Identify point charges or objects that can be modeled as point charges VISUALIZE: Use a pictoriai representation to establish a coordinate system, show the positions of the charges, show the force vectors on the charges, define cistances and ang es, and identfy what the problem is trying to find. This is the process of translating words to symbols SOLVE: The mathematical representation is based on Coulambs law . Show the directions of the forces repulsive for Eke charges, attractive far opposite charges- on the piciorial representation, . When possible, do graphical vector addition on the pictorial representation. Although not exac, it tells you the type of answer you should expect White each foroe vector in terms of its x and y components, and then add the components to find the net force. Use the pictorial representation to determine which components are positive and which are negative ASSESS: Check that your resut has the correct units, is reasonable, and answers the quesion. Review Learning Goal: To practice Problem-Solving Strategy 22.1 for electric torce problems. Two charged particlas, with chargos qand 4g, aro located on tha x axis soparated by a dstanco of 2.00 cm.A third charged particla, with chargo gs-g, is placod on tho x axis such that tha magnitude of the foroo hat chargo 1 oxorts on chargo 3 is oqual to the forco that chargo 2 axorts on chargo 3. Find the position of charge 3 when q-2.00 nC Part A Which of the following sketches represents a possible oonfiguration for this problem? 419392 41 43 92 42 41 42 s Enter the letter(s) indicating the correct graph(s) in alphabetical order. For example, if you think that A and B are correct, enter AB
Answer #1

Electrostatic force is given by

F = k*Q1*Q2/R^2

So force that charge 1 exerts on charge 3 will be

F13 = k*q1*q3/d1^2

where, d1 = distance between q1 and q3

Now force that charge 2 exerts on charge 3 will be

F23 = k*q2*q3/d2^2

where, d2 = distance between q2 and q3

Now given that

|F13| = |F23|

k*q1*q3/d1^2 = k*q2*q3/d2^2

given that

q1 = q & q2 = 4q & q3 = q, So

k*q*q/d1^2 = k*4q*q/d2^2

(d2/d1)^2 = 4

d2/d1 = sqrt 4 = 2

d2 = 2*d1

distance between q2 and q3 = twice of distance between q1 and q3

in other words d2 > d1

Now in above given configurations only A ad C follows above relation, So

Correct Answer: AC

Please Upvote.